OpenStudy (anonymous):

Find the standard form of the equation of the ellipse satisfying the given conditions. Minor axis with length 4 ; Foci at (-5,0), (5,0) Any help?

OpenStudy (tkhunny):

It's an ellipse with foci oriented parallel to the x-axis. It looks like this: \(\dfrac{(x-h)^{2}}{a^{2}} + \dfrac{(y-k)^{2}}{b^{2}} = 1\) Did you get that far?

OpenStudy (anonymous):

Yea, I know how to graph it when it's in standard form. I'm just having trouble getting in the form

OpenStudy (tkhunny):

Okay, where is the center of the ellipse. The foci give it away, right?